dontsaymyname:

Math Question for 702396

dontsaymyname:

1 attachment
dontsaymyname:

@702936

dontsaymyname:

The user doesn't know how to post questions, so I posted it for her, but she is the one who needs help.

702936:

yes i do

702936:

thank you

jhonyy9:

what method do you know to solve a system of equations ?

jhonyy9:

elimination or substitution what you need to use here ?

702936:

limitation

justus:

So for the first one we get this 2(9-2y)+y=12 18-3y=12 x=9-2*2 x=5 x=5,y=2 which box do you think matches it?

jhonyy9:

elimination ?

702936:

sorry yes

jhonyy9:

ok @justus do you wan help how use the elimination method to solve a system of equations ?

justus:

Sure this is an example.

jhonyy9:

in case of second there is +18y not minus

jhonyy9:

@justus

Imagine:

How to find the "X" Vaule

dontsaymyname:

Thank all of you for helping! She has the answer now! :)

justus:

Np and good luck !! :)))